find the slope 10x-3y=-60
Answer by RAY100(1637)

10x -3y = -60
.
rearrange
.
-3y = -10x -60
.
mult thru by (-1),,,, just to clarify
.
3y = 10x +60
.
divide all terms by 3
.
y = (10/3) x + 20
.
From the base form, y=mx +b, we can see that m= slope = 10/3
.
checking
.
spot a few points from origonal eqn,,,(0,20),,,(-6,0),,y2-y1/x2-x1 = 20/6=10/3,,,ok